# Time-Optimized Safe Navigation in Unstructured Environments through Learning Based Depth Completion

### v1

@@ 1 Introduction
Figure 1: A map of our environment to the left along with the flight path. Right is the color image and our generated camera depth
@@
Autonomous Micro Aerial Vehicles (MAVs) offer valuable solutions across a range of application scenarios such as agriculture [8], search-and-rescue [9], and infrastructure inspection [10]. To perform these maneuvers, the aerial vehicles have to locate the obstacles in real time while concurrently generating and executing safe and obstacle free trajectories.
@@
  • We develop a visual perception algorithm capable of generating dense and low-noise point clouds. Our method combines monocular, learning-based depth estimation with a traditional commercial off-the-shelf stereo matching solution. As a result, it produces depth estimates †with longer range, lower noise, and improved structural consistency compared to standard commercial solutions.
  • We propose the first real-time implementation in unknown environments for shortest path on the graph of convex sets (our path planning method). Our path planning method computation scales well with the size and complexity of maps outperforming other popular methods like Rapidly-exploring Random Tree (RRT).
  • We develop a unique time-optimized trajectory generation method with safety guarantees. We compare our trajectory generation method to MINCO [11] another time-optimized trajectory method used in multiple popular works [6, 2, 11] . Our method generates global trajectories more quickly while maintaining comparable optimality and ensuring safety.
  • We extensively validate our navigation system in both indoor (GPS denied) and outdoor environments for exploration tasks under varying lighting conditions and harsh winds of up to $15$\mathrm{m}\mathrm{p}\mathrm{h}$$, demonstrating reliable performance over areas of up to $60$\mathrm{m}$\times 40$\mathrm{m}$\times 3$\mathrm{m}$$.
@@ 2 Related Works
Mapping. There are two predominant sensors that are currently employed to generate environment depth onboard a quadrotor: cameras or LiDAR. Although other sensors exist, such as radar[12] or ultrasounds [13], they are far less prevalent on drones due to their size or low resolution. Most works using LiDARs [14, 7, 6, 15] are able to achieve strong mapping and localization performance. However, the LiDAR’s weight ranges from $250\leavevmode\nobreak\ $\mathrm{g}$$ to $1000\leavevmode\nobreak\ $\mathrm{g}$$ which limits their use on smaller aerial platforms. In contrast, stereo cameras offer a much lighter payload, typically less than $100\leavevmode\nobreak\ $\mathrm{g}$$. However small stereo camera pairs have low maximum depth sensing range, typically $3\leavevmode\nobreak\ $\mathrm{m}$-6\leavevmode\nobreak\ $\mathrm{m}$$ [2, 4, 5]. Learning based monocular depth estimation methods [16] offer an exciting new paradigm. Monocular depth estimation inferences depth based on the image’s structure by learning the mapping from images to depth on large datasets. Unfortunately, while the depth structure is relatively consistent, the scale of the depth is wildly off for objects outside the training set. One popular paradigm for addressing this problem is called depth completion, where a sparse metric depth supplements monocular information to create a dense structurally consistent metric depth. A supplemental sensor like a radar [17] can be fed into a larger neural network to infer the scale. However, this solution involves large datasets of unified sensor data to create this depth completion method therefore limiting their deployment. Unlike those methods, in the proposed work we develop an optimization-based scaling method using our sparse depth generated from our RGB-D camera similar to [3], but we employ inverse depth obtaining more accurate and consistent results. This method does not require retraining and is applicable to a variety of settings giving us great flexibility for deployment and switching the components in this system.
@@ 3.1 Mapping
In this section, we describe our depth generation method. We utilize an RGB-D camera that outputs temporally and spatially synchronized color and depth images. Our goal is to take the noisy depth image created by stereo and combine it with color information to create a less noisy and structurally consistent depth. First, we feed our color image ($W\times H\times 3$ dimensions) is fed to a monocular depth estimation network, DepthAnythingv2 [16], which generates an unscaled inverted depth $\mathbf{d}_{m}$ with units $\mathrm{m}\mathrm{m}^{-1}$. Monocular depth is an ill-posed problem. The structure created by the depth is relatively accurate which still makes it useful. However, the scale information is highly inaccurate. Conversely, stereo depth, $\mathbf{d}_{s}$ in $\mathrm{m}\mathrm{m}$, has a reliable scale but lacks structure and is filled with speckle noise and holes. Ideally, we want to fuse the monocular structure from the monocular depth with the scale from the stereo. To achieve this objective, we rescale our monocular depth with the following optimization problem
$\begin{split}\min_{\alpha_{2,1,0}}\sum_{j=0}^{W}\sum_{i=0}^{H}&\left(\frac{1}{\mathbf{d}_{s}[i,j]}-\alpha_{2}\mathbf{d}_{m}[i,j]^{2}-\alpha_{1}\mathbf{d}_{m}[i,j]-\alpha_{0}\right)^{2}\\ \mathbf{d}_{c}[i,j]&=\frac{1}{\alpha_{2}\mathbf{d}_{m}[i,j]^{2}+\alpha_{1}\mathbf{d}_{m}[i,j]+\alpha_{0}},\end{split}$ (1)
where $\mathbf{d}_{c}$ is the completed depth image in $\mathrm{m}\mathrm{m}$, $[\alpha_{2},\alpha_{1},\alpha_{0}]$ are scaling coefficients, and $i$ and $j$ are the index terms for the depth images. We choose a square approximation because it has been shown to be optimal in literature [3] for this problem. The scaling coefficients are solved using the least squares regression algorithm [36]. We perform our optimization on the $\mathrm{m}\mathrm{m}$^{-1}$ scale, as it consistently yields superior results compared to the $\mathrm{m}\mathrm{m}$ scale. In least squares regression, the largest values disproportionately influence the optimization outcome. When using the $\mathrm{m}\mathrm{m}$^{-1}$ scale, points closer to the camera have a greater impact on the optimization. Closer points to the camera benefit from higher stereo depth accuracy than distant points. In contrast, optimization on the $\mathrm{m}\mathrm{m}$ scale is dominated by distant points. As a result, the $\mathrm{m}\mathrm{m}$^{-1}$ scale is more suitable for achieving accurate depth optimization. The completed depth $\mathbf{d}_{c}$ and vehicle’s pose is fed to a mapping algorithm, Nvblox [37], to generate the occupancy map. After the occupancy is generated, we extract the free space and perform the cuboid decomposition described in the following.
@@ 3.2 Cuboid Decomposition and Path Selection
To generate an obstacle-free, dynamically feasible trajectory, our approach start by generating a set of Safe Flight Corridors (SFC) connecting the start and goal. Most existing methods [11, 19, 20] first compute a obstacle free path and then create SFC around the path. Our method instead reverses this procedure converting the occupancy map to a graph of SFC, and then finding the subset of SFC that links start and goal. We call our clustering approach to turn the occupancy map into a graph of safe flight corridors through cuboid decomposition. The advantage of our approach is that the clustering can be performed at the map generation step in parallel to trajectory planning before a goal is given. Additionally, once the clustering is complete, solving the subset of safe flight corridors is faster scaling with the number of SFC in the graph rather than size of the map.
@@
  Initialization: Mark all free space as uncovered.
@@
In the literature, cuboid decomposition can be viewed as a covering problem, where the goal is to cover the free space with boxes, as described in [38]. Our proposed cuboid decomposition algorithm is illustrated in Fig. 3. The procedure involves the following steps described in our Algorithm  1. In general, we first repeatedly create cuboids to cover the free space. Subsequently, we construct the edges of our graph by checking if each pair is connected. Each vertex of this graph represents a cuboid of free space and each edge represents an adjacency. This graph-based representation is updated repeatedly, independently of path planning. To extend this decomposition into $2.5$D, we divide the space into several $2$D layers and repeat the process for each layer.
Once we have a graph representation of the free space, we solve for our path. Rather than an explicit path, we are selecting a set of SFC that connect our start and goal positions. We implement a modified A* algorithm to achieve our goal. The edge weight of each corridor is $1$. The lower bound heuristic is the minimum distance between a point in the cuboid node and the goal, assuming that no obstacles exist. As the number of cuboids is significantly lower than the nodes in an occupancy grid, A* can be solved in less than $1\leavevmode\nobreak\ $\mathrm{m}\mathrm{s}$$ while providing a near-optimal set of safe flight corridors as seen in Fig. 4c. Solving for a true optimal set of safe flight corridors is a NP-Hard problem [28]. Conversely, this algorithm’s objective is to minimize the number of SFC which in turn corresponds to minimizing the number of optimization variables for our trajectory generation.
@@ 3.3 Bernstein Time-Optimized Trajectory Planning
We solve for a trajectory that connects a start condition of position, velocity, and acceleration $[\mathbf{x}_{0}^{\top},\mathbf{v}_{0}^{\top},\mathbf{a}_{0}^{\top}]^{\top}$ with a goal position $\mathbf{x}_{f}$ given a set of safe flight corridors. We introduce the formalism for a single-dimensional case. To extend this to multiple dimensions, it is sufficient to repeat the constraints for each dimension. We formulate our trajectory using Bernstein polynomial splines. Each spline is assigned to a single safe flight corridor. First, we define each spline as follows
$\begin{split}B(t)&=\begin{cases}\text{$B_{0}\left(t\right)$}&\quad\text{if $t\in\left[0,t_{0}\right]$}\\ \text{$B_{1}\left(t-t_{0}\right)$}&\quad\text{if $t\in\left[t_{0},t_{0}+t_{1}\right]$}\\ \hskip 10.0pt\vdots\\ \text{$B_{f}\left(t-\sum\limits_{i=1}^{f-1}t_{i}\right)$}&\quad\text{if $t\in\left[\sum\limits_{i=1}^{f-1}t_{i},\sum\limits_{i=1}^{f}t_{i}\right]$}\\ \end{cases},\\ \end{split}$ (2)
@@
$\begin{split}B_{i}(t)&=\sum_{k=0}^{5}c^{k}_{i}b^{5}_{k}\left(\frac{t}{t_{i}}\right),\\ \end{split}$ (3)
where $c^{k}_{i}$ is the coefficient of the polynomial, $k$ represents the coefficient number and $i$ is the spline number. In literature, an alternative name for the coefficient is control points when describing Bernstein polynomials. $b^{5}_{k}(t)$ represents the basis function defined in [34] for order $5$ and 1-D polynomial.
However instead of structuring the problem for coefficients, we structure our optimization as solving for a set of waypoints $\mathbf{w}=[\mathbf{w}_{0},\mathbf{w}_{1},...,\mathbf{w}_{f}]^{\top}$ and traversal times $\mathbf{t}=[t_{0},t_{1},...t_{f}]^{\top}$ We define a waypoints as a $\mathbf{w}_{i}\in\mathbb{R}^{3\times 1}$ with elements $\mathbf{w}_{i}=\begin{bmatrix}B_{i}(0)&B_{i}^{\prime}(0)&B_{i}^{\prime\prime}(0)\end{bmatrix}^{\top}$ where $B^{\prime}=\frac{dB(t)}{dt}$ is the velocity, and $B^{\prime\prime}=\frac{d^{2}B(t)}{dt^{2}}$ is acceleration. Each spline is completely defined by the start and end waypoint on the curve. This relationship between coefficients and waypoints is defined as
$\begin{bmatrix}c^{0}_{i}\\ c^{1}_{i}\\ \vdots\\ c^{4}_{i}\\ c^{5}_{i}\\ \end{bmatrix}=\begin{bmatrix}1&0&0&0&0&0\\ 1&\frac{1}{5}t_{i}&0&0&0&0\\ 1&\frac{2}{5}t_{i}&\frac{1}{20}(t_{i})^{2}&0&0&0\\ 0&0&0&1&-\frac{2}{5}t_{i}&\frac{1}{20}(t_{i})^{2}\\ 0&0&0&1&-\frac{1}{5}t_{i}&0\\ 0&0&0&1&0&0\\ \end{bmatrix}\begin{bmatrix}\mathbf{w}_{i}\\ \mathbf{w}_{i+1}\\ \end{bmatrix},$ (4)
Building on this structure, we formulate the trajectory generation problem with the following optimization
$\min_{\mathbf{w}_{[0,...,f]},\mathbf{t}_{[0,...,f]}}$ $\leavevmode\nobreak\ \sum^{f}_{i=0}t_{i},$ (5)
@@
Structuring the problem using waypoints instead of coefficients improves our optimization in three ways. First, optimizing on waypoints reduces the number of optimization variables compared to coefficients when increasing the number of splines. For example, with $f$ number of splines of order $5$, the number of optimization variables using waypoints is $3(f+1)$ while using coefficients it requires $6f$. Next, the gradients are much better conditioned on functions like $t_{i}^{2}$ and $t_{i}$ as opposed to optimizing over coefficients which involve the inverse, $\frac{1}{t_{i}}$ or $\frac{1}{{t_{i}}^{2}}$. Finally, it is easier to initialize our optimizer because waypoints have a better physical interpretation with respect to coefficients.
Subsequently, we define our safety constraints. The safe flight corridors are defined by their position constraints lower $\mathbf{l}=[l_{0},l_{1},...l_{f}]^{\top}$ and upper $\mathbf{u}=[u_{0},u_{1},...u_{f}]^{\top}$ for each corridor. With these bounds, we formulate our safety constraints from eq. (5), $\mathbf{y}$, $\mathbf{G},\mathbf{z}$, based on the convex hull property of Bernstein polynomials [34]. The convex hull property states bounds on the coefficients are equivalent to bounds on the full trajectory defined formally as
$\begin{split}l_{i}\leq\mathbf{A}_{i}\begin{bmatrix}\mathbf{w}_{i}\\ \mathbf{w}_{i+1}\\ \end{bmatrix}\leq u_{i}&\implies l_{i}\leq B_{i}(t)\leq u_{i}\hskip 5.0pt,\\ -\mathbf{v}_{m}\leq\mathbf{D}^{(1)}_{i}\mathbf{A}_{i}\begin{bmatrix}\mathbf{w}_{i}\\ \mathbf{w}_{i+1}\\ \end{bmatrix}\leq\mathbf{v}_{m}&\implies-\mathbf{v}_{m}\leq B^{\prime}_{i}(t)\leq\mathbf{v}_{m},\\ \underbrace{-\mathbf{a}_{m}\leq\mathbf{D}_{i}^{(2!)}\mathbf{A}_{i}\begin{bmatrix}\mathbf{w}_{i}\\ \mathbf{w}_{i+1}\\ \end{bmatrix}\leq\mathbf{a}_{m}}_{\text{Optimization Constraint}}&\implies\underbrace{-\mathbf{a}_{m}\leq B^{\prime\prime}_{i}(t)\leq\mathbf{a}_{m}}_{\text{Physical Result}},\\ \end{split}$ (6)
@@
$\begin{split}\mathbf{D}^{(m)}_{i}&=\frac{6-m}{t_{i}}\begin{bmatrix}-1&1&0&0&\ldots\\ 0&-1&1&0&\ldots\\ &&\ddots&\ddots&\\ \end{bmatrix},\\ \mathbf{D}^{(m!)}_{i}&=\mathbf{D}^{(1)}_{i}\mathbf{D}^{(2)}_{i}\ldots\displaystyle\mathbf{D}^{(m)}_{i}.\end{split}$ (7)
We then generate the initial trajectory $\mathbf{w}_{\text{init}}$. This trajectory is a straight line trajectory between the start point and the center of the intersections of each safe flight corridors visualized in Fig. 4d or mathematically defined as
$\begin{split}\mathbf{x}_{i}&=\frac{1}{2}(\max[l_{i},l_{i+1}]+\min[u_{i},u_{i+1}]),\\ \mathbf{w}_{i}&=\begin{bmatrix}\mathbf{x}_{i}^{\top}&0&0\\ \end{bmatrix}^{\top},\hskip 10.0pt1\leq i\leq f-1,\\ \mathbf{w}_{\text{init}}&=\begin{bmatrix}\mathbf{w}_{0}^{\top}&\mathbf{w}_{1}^{\top}&\ldots&\mathbf{w}_{i}^{\top}&\ldots&\mathbf{w}_{f}^{\top}\\ \end{bmatrix}^{\top},\\ \mathbf{t}_{\text{init}}&=\begin{bmatrix}\frac{||\mathbf{x}_{1}-\mathbf{x_{0}}||}{||\mathbf{v}_{m}||}&\frac{||\mathbf{x}_{2}-\mathbf{x_{1}}||}{||\mathbf{v}_{m}||}&\ldots&\end{bmatrix}^{\top}.\end{split}$ (8)
@@
$\psi_{d}(t)=\begin{cases}\psi_{c}&\quad\text{if $||v_{y}||,||v_{x}||<1e-1$}\\ \psi_{c}-\psi_{m}&\quad\text{if $\arctan(v_{y},v_{x})-\psi_{c}<-\psi_{m}$}\\ \psi_{c}+\psi_{m}&\quad\text{if $\arctan(v_{y},v_{x})-\psi_{c}>\psi_{m}$}\\ \arctan(v_{y},v_{x})&\quad\text{else}\end{cases}$ (9)
@@
To handle newly observed obstacles, we implement a repeated replanning procedure. First, we estimate if our currently planned trajectory intersects with the updated map’s obstacles within the next $10\leavevmode\nobreak\ $\mathrm{s}$$. If no collision is detected, then we continue with this flight plan. Otherwise, we plan a new trajectory with the updated map to our goal. In the case where no new trajectory can be solved, we stop the quadrotor (for example the goal is in a wall). This check of deciding whether to continue, stop, or replan is repeated at a rate of $20\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$.
@@
Figure 8: Depth Comparison. D455 Depth in blue is the stereo depth generated directly from the camera. Our Depth is the depth completion method described in our methodology.

### v2

@@ 1 Introduction
Figure 1: A map of our environment on the top. Bottom right is the color image, and bottom Left is our depth completion.
@@
Autonomous Micro Aerial Vehicles (MAVs) offer valuable solutions across a range of application scenarios such as agriculture [8], search-and-rescue [9], and infrastructure inspection [10]. To services these applications, the aerial vehicles have to locate the obstacles in real time while concurrently generating and executing safe and obstacle free trajectories.
@@
  • We develop a visual perception algorithm capable of generating dense and low-noise depth based on depth completion from learning based depth and classical computer vision.
  • We propose the first real-time implementation for shortest path on the graph of convex sets (our path planning method) in unknown environments.
  • We develop a unique time-optimized trajectory generation method with safety guarantees along with a mathematical derivation of our polynomial properties. Our method generates global trajectories more quickly while ensuring safety.
  • We integrate the proposed algorithms into a complete software stack for real time onboard aerial autonomous navigation in unstructured environments that relies solely on lightweight onboard sensors, namely cameras and an IMU.
  • We extensively validate our navigation system in both indoor (GPS denied) and outdoor environments for exploration tasks under varying lighting conditions and harsh winds of up to $15$\mathrm{m}\mathrm{p}\mathrm{h}$\$, demonstrating reliable performance over areas of up to $60$\mathrm{m}$\times 40$\mathrm{m}$\times 3$\mathrm{m}$\$.
@@ 2 Related Works
Mapping. There are two predominant sensors that are currently employed to generate environment depth onboard a quadrotor: cameras or LiDAR. Although other sensors exist, such as radar[11] or ultrasounds [12], they are far less prevalent on drones due to their size or low resolution. Most works using LiDARs [13, 7, 6, 14] are able to achieve strong mapping and localization performance. However, the LiDAR’s weight ranges from $250\penalty 10000\ $\mathrm{g}$\$ to $1000\penalty 10000\ $\mathrm{g}$\$ which limits their use on smaller aerial platforms. In contrast, stereo cameras offer a much lighter payload, typically less than $100\penalty 10000\ $\mathrm{g}$\$. However small stereo camera pairs have low maximum depth sensing range, typically $3\penalty 10000\ $\mathrm{m}$-6\penalty 10000\ $\mathrm{m}$\$ [2, 4, 5]. Learning based monocular depth estimation methods [15] offer an exciting new paradigm. Monocular depth estimation inferences depth based on the image’s structure by learning the mapping from images to depth on large datasets. Unfortunately, while the depth structure is relatively consistent, the scale of the depth is wildly off for objects outside the training set. One popular paradigm for addressing this problem is called depth completion, where a sparse metric depth supplements monocular information to create a dense structurally consistent metric depth. A supplemental sensor like a radar [16] can be fed into a larger neural network to infer the scale. However, this solution involves large datasets of unified sensor data to create this depth completion method therefore limiting their deployment. Unlike those methods, in the proposed work we develop an optimization-based scaling method using our sparse depth generated from our RGB-D camera similar to [3], but we employ inverse depth obtaining more accurate and consistent results. This method does not require retraining and is applicable to a variety of settings giving us great flexibility for deployment and switching the components in this system.
@@ 3 Methodology
Estimated Value True Depth Depth Error Disparity Error
$1$\mathrm{m}$,2$\mathrm{m}$,5$\mathrm{m}$\$] $1$\mathrm{m}$,2$\mathrm{m}$,10$\mathrm{m}$\$] $5$\mathrm{m}$\$ $\mathbf{0.1$\mathrm{p}\mathrm{x}$}$
$2$\mathrm{m}$,4$\mathrm{m}$,10$\mathrm{m}$\$] $1$\mathrm{m}$,2$\mathrm{m}$,10$\mathrm{m}$\$] $\mathbf{2.23$\mathrm{m}$}$ $0.55$\mathrm{p}\mathrm{x}$\$
Table 2: Depth error versus disparity error examples. We assume 33 objects and compare them to a ground truth depth. We see that depth error focuses on amplifies the error on the far objects. Disparity on the other hand amplifies the error on close objects. We assume a baseline of $1$\mathrm{m}$\$ and focal length of $1$\mathrm{m}$\$.
@@ 3.1 Mapping
In this section, we describe our depth generation method. Without accurate mapping, the following sections on planning and control are unable to be implemented in unknown environments. By solving this problem, we are able to ensure an accurate map to leverage our planning and control for safe navigation. We utilize an RGB-D camera that outputs temporally and spatially synchronized color and stereo disparity images. Our goal is to take the noisy disparity image created by stereo and combine it with color information to create a less noisy and structurally consistent disparity which will then be converted to depth. We prefer to do our calculations in this disparity or inverted depth scale for our application of navigation and image based sensors rather than normal depth scale. This allows us to emphasize error on the close objects where image based sensors are more accurate and is more relevant for planning.
We will use the following example to elucidate our point. Imagine our image has $3$ main objects with a true distance of $1$\mathrm{m}$,2$\mathrm{m}$,10$\mathrm{m}$\$]. Then let’s say we have $2$ methods that give separate distance estimates. Method A estimates $1$\mathrm{m}$,2$\mathrm{m}$,5$\mathrm{m}$\$]. Method B estimates $2$\mathrm{m}$,4$\mathrm{m}$,10$\mathrm{m}$\$]. We see various error results in Table. 2. Depth error is computing by directly using the RMSE. Disparity error is inverting the estimated value and true depth and then calculating the RMSE. We see that choosing disparity error amplifies the error for close range objects over the long range objects done in the depth error. From a planning perspective, disparity error is superior because accurately representing the distance of close objects is more important than the far object. This is especially aggravated since our system is continuously replanning, and we will have less time to react to close objects as opposed to far objects when we replan.
To perform our depth completion, we do the following steps. First, we feed our color image ($W\times H\times 3$ dimensions) is fed to a monocular depth estimation network, DepthAnythingv2 [15], which generates an unscaled estimated disparity image $\mathbf{d}_{m}$ with units $\mathrm{p}\mathrm{x}$. Monocular depth is an ill-posed problem. The structure created by the neural network is good however the scale information is highly inaccurate. Conversely, stereo disparity, $\mathbf{d}_{s}$ in pixels or $\mathrm{p}\mathrm{x}$, has a reliable scale but lacks structure and is filled with speckle noise and holes. This is visible in Fig. 2 where the stereo depth and from it disparity has the several large holes in the building structures. Ideally, we want to fuse the structure from the monocular depth with the scale from the stereo. To achieve this objective, we rescale our monocular depth with the following optimization problem
$\begin{split}\mathbf{d}_{c}[i,j]&=\alpha_{2}\mathbf{d}_{m}[i,j]^{2}+\alpha_{1}\mathbf{d}_{m}[i,j]+\alpha_{0},\\ \min_{\alpha_{2,1,0}}\sum_{j=0}^{W}\sum_{i=0}^{H}&\left(\mathbf{d}_{s}[i,j]-\mathbf{d}_{c}[i,j]\right)^{2}.\end{split}$ (1)
[…]
@@ 3.2 Cuboid Decomposition and Path Selection
To generate an obstacle-free, dynamically feasible trajectory, our approach start by generating a set of Safe Flight Corridors (SFC) to describe the map. Most existing methods [29, 18, 19] first compute a obstacle free path and then create SFC around the path. Our method instead reverses this procedure converting the occupancy map to a graph of SFC, and then finding the subset of SFC that links start and goal. We call our approach to turn the occupancy map into a graph of safe flight corridors cuboid decomposition. The advantage of our approach is that the clustering can be performed at the map generation step in parallel to trajectory planning before a goal is given. Additionally, once the clustering is complete, solving the subset of safe flight corridors to connect the starting pose and goal is orders of magnitude faster.
@@
Initialization: Expand the obstacles by the control error and vehicle’s radius. Mark all free space as uncovered.
@@
Figure 4: Spline Definition: $\mathbf{w}_{i}$ the waypoint number. $\mathbf{B}_{i}$ Bernstein spline between each waypoints.
In the literature, cuboid decomposition can be viewed as a covering problem, where the goal is to cover the free space with boxes, as described in [38]. Our proposed cuboid decomposition algorithm is illustrated in Fig. 3. The procedure involves the following steps described in our Algorithm 1. We initialize the space by expanding the obstacles by the average control error, localization error, and vehicle radius. Then we set all free space to uncovered. After initialization, we repeatedly create cuboids to cover the free space. Subsequently, we construct the edges of our graph by checking if each pair of cuboids share points. As a result, each vertex of this graph represents a cuboid of free space and each edge represents an adjacency. This graph-based representation is updated repeatedly, independently of path planning. To extend this decomposition into 2.52.5D, we divide the space into several 22D layers and repeat the process for each layer. This process is done by declaring a minimum and maximum flying height for the vehicle. We then divide this minimum and maximum height equally among the $3$ layers along with some buffer height.
Once we have a graph representation of the free space, we solve for our path. Rather than an explicit path, we are selecting a set of SFC that connect our start and goal positions. We implement a modified A* algorithm to achieve our goal. The edge weight of each corridor is $1$. The lower bound heuristic is the minimum distance between a point in the cuboid node and the goal, assuming that no obstacles exist. We include in our appendix why other edge weight metrics such as centers of the convex sets or minimum distance between sets is less optimal than a constant edge weight. As the number of cuboids is significantly lower than the nodes in an occupancy grid, A* can be solved in a fraction of time while providing a near-optimal set of safe flight corridors as seen in Fig. 5c. Solving for a true optimal set of safe flight corridors is a NP-Hard problem [27]. Conversely, this algorithm’s objective is to minimize the number of SFC which in turn corresponds to minimizing the number of optimization variables for our trajectory generation.
@@ 3.3 Bernstein Time-Optimized Trajectory Planning
We solve for a trajectory that connects a start condition of position, velocity, and acceleration $\mathbf{x}_{0}^{\top},\mathbf{v}_{0}^{\top},\mathbf{a}_{0}^{\top}]^{\top}$ with a goal position $\mathbf{x}_{f}$ given a set of safe flight corridors. We introduce the formalism for a single-dimensional case. To extend this to multiple dimensions, it is sufficient to repeat the constraints for each dimension. We formulate our trajectory using a fifth order Bernstein polynomial splines. Each spline is assigned to a single safe flight corridor. First, we define each spline as follows
[…]
@@
where $c_{k,i}$ is the coefficient of the polynomial, $k$ represents the coefficient number and $i$ is the spline number. In literature, an alternative name for the coefficient is control points when describing Bernstein polynomials. $b_{k,5}(t)$ represents the basis function defined in [34] for order $5$ and 1-D polynomial. We can see in fig. 4 a visualization of our spline.
However instead of the optimization problem to solve for coefficients, we structure our optimization as solving for a set of waypoints $\mathbf{w}=[\mathbf{w}_{0},\mathbf{w}_{1},...,\mathbf{w}_{f}]^{\top}$ and traversal times $\mathbf{t}=[t_{0},t_{1},...t_{f-1}]^{\top}$ . We define a waypoints as a $\mathbf{w}_{i}\in\mathbb{R}^{3\times 1}$ with elements $\mathbf{w}_{i}=\begin{bmatrix}B_{i}(0)&B_{i}^{\prime}(0)&B_{i}^{\prime\prime}(0)\end{bmatrix}^{\top}$ where $B^{\prime}=\frac{dB(t)}{dt}$ is the velocity, and $B^{\prime\prime}=\frac{d^{2}B(t)}{dt^{2}}$ is acceleration. These waypoints are visualized in fig. 4. In our formulation, we have $f$ splines and $f+1$ waypoints describing these splines. Each spline is completely defined by the start and end waypoint on the curve. This relationship between coefficients and waypoints is defined as
$\begin{bmatrix}c_{0,i}\\ c_{1,i}\\ \vdots\\ c_{4,i}\\ c_{5,i}\\ \end{bmatrix}=\begin{bmatrix}1&0&0&0&0&0\\ 1&\frac{1}{5}t_{i}&0&0&0&0\\ 1&\frac{2}{5}t_{i}&\frac{1}{20}(t_{i})^{2}&0&0&0\\ 0&0&0&1&-\frac{2}{5}t_{i}&\frac{1}{20}(t_{i})^{2}\\ 0&0&0&1&-\frac{1}{5}t_{i}&0\\ 0&0&0&1&0&0\\ \end{bmatrix}\begin{bmatrix}\mathbf{w}_{i}\\ \mathbf{w}_{i+1}\\ \end{bmatrix}$, (4)
The derivation of eq. 4 is included in our appendix. Building on this structure, we formulate the trajectory generation problem with the following optimization
$\min_{\mathbf{w}_{[0,...,f]},\mathbf{t}_{[0,...,f-1]}}$ $\penalty 10000\ \sum^{f-1}_{i=0}t_{i},$ (5)
$\textrm{s.t.}\quad\mathbf{w}_{0}$ $=\begin{bmatrix}\mathbf{x}_{0}^{\top}&\mathbf{v}_{0}^{\top}&\mathbf{a}_{0}^{\top}\end{bmatrix}^{\top}$,
@@
Structuring the problem using waypoints instead of coefficients improves our optimization in three ways. First, optimizing on waypoints reduces the number of optimization variables compared to coefficients when increasing the number of splines. For example, with $f$ number of splines with $6$ coefficients and $1$ time variable per spline, we would have $7f$ optimization variables. Using our method, the number of optimization variable is 3(f+1)+f63(f+1)+f-6 or 4f34f-3 optimization variables. We have $f+1$ waypoints, $f$ time variables, and since the start and end waypoints are set to values we can replace them with constants resulting in 6-6 optimization variables. Next, the gradients are much better conditioned on functions like $t_{i}^{2}$ and $t_{i}$ as opposed to optimizing over coefficients which involve the inverse, $\frac{1}{t_{i}}$ or $\frac{1}{{t_{i}}^{2}}$ like in eq. 23. Finally, it is easier to initialize our optimizer because waypoints have a better interpetability than coefficients.
Subsequently, we define our safety constraints. The safe flight corridors are defined by their position constraints lower $\mathbf{l}=[l_{0},l_{1},...l_{f-1}]^{\top}$ and upper $\mathbf{u}=[u_{0},u_{1},...u_{f-1}]^{\top}$ for each corridor. With these bounds, we formulate our safety constraints from eq. (5), $\mathbf{y}$, $\mathbf{G},\mathbf{z}$, based on the convex hull property of Bernstein polynomials [34]. We prove this property in the appendix of this work as well. The convex hull property states bounds on the coefficients are equivalent to bounds on the full trajectory defined formally as
[…]
@@
[…]
This matrix derivation in eq. 7 is also included in our appendix. We then generate the initial trajectory $\mathbf{w}_{\text{init}}$. This trajectory is a straight line trajectory between the start point and the center of the intersections of each safe flight corridors visualized in Fig. 5d or mathematically defined as
[…]
@@
[…]
@@
To handle newly observed obstacles, we implement a repeated replanning procedure. First, we estimate if our currently planned trajectory intersects with the updated map’s obstacles within the next $10\penalty 10000\ $\mathrm{s}$\$. If no collision is detected, then we continue with this flight plan. Otherwise, we plan a new trajectory with the updated map to our goal. In the case where no new trajectory can be solved, we stop the quadrotor (for example the goal is in a wall). This check of deciding whether to continue, stop, or replan is repeated at a rate of $20\penalty 10000\ $\mathrm{H}\mathrm{z}$\$.
@@
Figure 9: Depth Comparison. D455 Depth in blue is the stereo depth generated directly from the Intel Realsense D455 Depth Camera. This is a classical computer vision algorithm similar to block matching. Our Depth is the depth completion method described in our methodology. We see performing our method already gives superior results than the consumer grade depth camera.